# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 12.7 Malatya’s Journey to Global Leadership: Implementing Best Practices
The other two crucial factors are the capacity to conduct larger exchanges and, more importantly, the inclusion of compatible pairs in the system. These are the areas in which the Malatya team truly excelled.
When our collaboration began in 2019, the Malatya team had already demonstrated the ability to perform five simultaneous liver transplants. Although the Liver Transplantation Institute at İnönü University was equipped to handle this capacity, they had the personnel to support up to seven. Following the world’s first 4-way exchange in July 2022, the team added two operating rooms to the existing ten at the Institute—all located on a single floor—enabling 6-way exchanges within its premises. By December 2024, they had utilized their expanded capacity for 7-way exchanges twice, using additional operating rooms within the larger hospital complex.
However, this impressive logistical capacity was not the primary driver of their extraordinary performance after the program’s pilot phase. Rather, it was the integration of “biological” incentive mechanisms, encouraging the inclusion of compatible pairs, that played the most significant role in their success.
Following the highly successful media event at the end of July 2023, Yılmaz inquired about additional policies that could enhance the efficacy of the BBS-LPE System. Reminding him that our 4-way exchange in July 2022 had been made possible by a blood-type O donor paired with a blood-type A patient, we asked what other strategies might encourage such pairs to participate in liver exchange. He explained that there are several reasons why blood-type compatible pairs may benefit from participating in liver exchange.
@@
These scenarios have been utilized by the BBS-LPE System since its deployment. However, three additional factors could further enhance its efficacy.
Firstly, liver transplants using blood-type identical grafts are generally better for patients than those using blood-type compatible but non-identical grafts (Gordon et al., 1986; Koukoutsis et al., 2007). Although the patient may not have antibodies against the transplanted liver, the liver itself can temporarily carry antibodies against the patient’s body. This phenomenon, known as passenger lymphocyte syndrome, often leads to hemolytic anemia in recipients of blood-type non-identical grafts. While this condition typically resolves on its own, it can occasionally result in severe complications, including death. In fact, in medical practice, transplanting a blood-type compatible but non-identical graft is referred to as minor ABO incompatibility (Bruijn et al., 2017). This creates an additional “biological” incentive to enhance the efficacy of the BBS-LPE System.
@@
In addition to integrating these “biological” incentive mechanisms into the BBS-LPE System, the Malatya team has also allowed for “altruistic” participation from compatible pairs, even when there is no clear benefit for the pair, provided that the patient receives an equally suitable graft without delay.
The adoption of these policies since August 2023 has been the primary driver behind the exceptional outcomes of the BBS-LPE program following its pilot phase. Our hope is that these practices will evolve into established standards in liver exchange, enhancing its global efficacy by orders of magnitude.
@@ 13 Contributions, Insights and Directions on Behavioral Market Design
While I have primarily relied on theory as the formal apparatus for my pursuits in minimalist market design, many applications have revealed intriguing insights, questions, and puzzles within behavioral economics. In this section, I explore some of the most significant ones.
@@ 13.1 Leveling the Playing Field in School Choice
To build a compelling case against this mechanism in favor of the Deferred Acceptance (DA) or Top Trading Cycles (TTC) mechanisms, Yan Chen and I conducted a laboratory experiment involving all three (Chen and Sönmez, 2006).138138138Building on Chen and Sönmez (2006), a substantial body of literature has employed lab experiments to study preference manipulation in school choice mechanisms. For a comprehensive review of this literature and other behavioral studies in education market design, see Rees-Jones and Shorrer (2023). For a broader review of behavioral studies in market design, see Chen et al. (2021). In our experimental setting, both DA and TTC outperformed the Boston mechanism.
@@
In our experiment, we deliberately withheld this key information, not disclosing the strategy-proofness of DA and TTC to participants. Our goal was to explore whether these mechanisms could still outperform Boston without any external guidance. Both mechanisms did, with DA performing best. However, a substantial number of participants still submitted non-truthful preferences under both DA and TTC—albeit less frequently than under the Boston mechanism.
Encouraged by these experimental results, we grew confident that replacing the Boston mechanism with either DA or TTC could yield even greater benefits in real-world settings, especially when families are properly informed and guided.
@@
Fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. The set of Nash equilibrium outcomes of this game is equal to the set of outcomes that satisfy individual rationality, non-wastefulness, and no justified envy.139139139Equivalently, by Proposition 4, the set of Nash equilibrium outcomes of this game is equal to the set of stable outcomes of the induced many-to-one matching market.
@@
Fix a school choice problem. The outcome of the DA mechanism under truthful dominant strategies either Pareto dominates or is equal to any given Nash equilibrium outcome of the preference revelation game induced by the Boston mechanism.
During our first meeting with Boston Public Schools (BPS) officials in October 2003, the experimental evidence from Chen and Sönmez (2006) and the theoretical results from Ergin and Sönmez (2006) were instrumental in helping to persuade BPS officials that it might be in the city’s best interest to consider replacing the Boston mechanism with either the DA or TTC mechanism (see Section 6.10).
Even though the reform process took close to two years for due diligence, it was clear from the beginning that both BPS leadership and the members of the Task Force formed to evaluate the city’s assignment process were supportive of a reform that would increase the transparency of the student assignment mechanism. While the decision between DA and TTC remained uncertain for some time, it was clear that there was little interest in retaining the Boston mechanism. Everything was progressing just as I had hoped.
In his memo to the School Committee on May 25, 2005, Superintendent Thomas Payzant wrote (Abdulkadi̇roğlu et al., 2006):
To my knowledge, this was the first time strategy-proofness—a technical axiom from mechanism design—had ever been invoked by a policymaker in public deliberations to advocate a position. Moreover, rather than its main interpretation in mechanism design, Payzant framed strategy-proofness as an equity principle.
@@
Subsequent analysis of school choice data from BPS under the Boston mechanism confirmed the widespread presence of both highly strategic and relatively naive preference submissions (Abdulkadi̇roğlu et al., 2006). For example, among the latter group, 19% of families listed two consistently oversubscribed (i.e., popular) schools as their top two choices, leading to 27% of these children being unassigned.
@@
Partitioning the students into strategic and sincere groups, fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. There exists a Nash equilibrium outcome in this game that is weakly preferred by each strategic student to the outcome of the DA mechanism in truthful dominant strategies.140140140With the BPS data used in Pathak and Sönmez (2008), the Nash equilibrium outcome of the preference revelation game induced by the Boston mechanism is virtually unique under all partitions of students into strategic and sincere.
@@
We now turn to a sharply contrasting case study from Seattle, where the school assignment process remained opaque for nearly a decade, possibly to serve objectives of decision makers that, by my definition, do not qualify as “legitimate.”
@@ 13.2 Opting for Obscurity: The Curious Case of Seattle School Choice
Up until then, all of us in the market design community believed that no jurisdiction worldwide had used the DA mechanism for student assignment prior to the publication of Abdulkadiroğlu and Sönmez (2003). Thanks largely to MacGregor’s efforts, we discovered that Seattle had transitioned from the Boston mechanism to the DA mechanism in 1999, through what was known as the Barnhart-Waldman (BW) amendment. This amendment, named after two school board members familiar with the U.S. entry-level job market’s use of DA, had led to the change. However, the switch was poorly communicated, if at all, resulting in widespread confusion among both the families participating in the system and the officials responsible for explaining it.
In Pathak and Sönmez (2013), we identified several indicators that the BW amendment was not well understood, even though a strategy-proof mechanism should have enabled the school district to advise families that submitting truthful preferences was a safe strategy. Confusion about the BW amendment even surfaced during a 2001 deposition of the school board president in a court challenge to Seattle’s school choice plan.141141141Page 58 of the U.S. Court of Appeals for the Ninth Circuit, No. 01–35450, Parents Involved in Community Schools vs. Seattle School District No.1, 2001.
@@
This also raises another issue. In Section 13.1, we have seen that even with full transparency, some participants in Boston utilized the available information better than others. In Seattle, judging from the inconsistent advice provided by officials, such as school principals, it is likely that while some families understood that truthful preference revelation was in their best interest, others did not. These families surely gained an even greater advantage than members of the WZPG in Boston compared to less informed families. In general, inconsistent communication of various aspects of an institution can even be exploited by decision makers to benefit some groups at the expense of others.
@@
In June 2009, the Seattle School Board proposed a new student assignment plan to transition from the DA mechanism back to the Boston mechanism. This proposal faced significant opposition from families. Many of the concerns raised, noted in Walkup (2009), echoed those discussed in Abdulkadiroğlu and Sönmez (2003) regarding the limitations of the Boston mechanism.
@@
Perplexed by this reversal—the only instance we were aware of at the time where a jurisdiction had replaced a school choice mechanism with a more manipulable one—we corresponded with several committee members involved in the decision. In Pathak and Sönmez (2013), we reported the only plausible explanation we encountered: that the BW amendment (i.e., the DA mechanism) would increase transportation costs by encouraging student mobility, as parents could freely express their true preferences without risking a seat at their neighborhood schools. If this was indeed the case, it likely became a concern only after the strategy-proofness of the city’s DA mechanism was brought to public attention, thanks to MacGregor’s efforts.
If reducing transportation costs was indeed a factor in this reform, it would exemplify an objective that does not meet my definition of “legitimate.” Rather than directly addressing school choice, were the decision makers attempting to limit the mobility it introduced indirectly by reverting to the Boston mechanism? This rationale would also be consistent with the decision to keep the strategy-proofness of the school choice mechanism opaque for nearly a decade—from 1999 to 2008.
@@
Beyond Seattle’s experience, could the Boston mechanism be a preferred mechanism for decision makers who favor neighborhood-based assignment policies but avoid directly advocating for them due to political costs?
@@ 13.3 Nudging School Choices: Taiwan’s Deduction Mechanism
Despite these relatively modest deductions, it is evident that—partly due to the mechanism’s explicit penalties for lower-ranked choices—TD mechanisms other than DA are all susceptible to preference manipulation. Thus, of all TD mechanisms, only DA satisfies the axiom of strategy-proofness. Similarly, none of the TD mechanisms, except for DA, satisfy the axiom of no justified envy.142142142It’s important to note that the school priority rankings induced by the effective scores under the TD mechanism are artifacts of the mechanism, not inherent to the problem. Therefore, no justified envy is defined based on the original student scores from the CAE.
@@
In response, Education Minister Chiang Wei-ling and K-12 Education Administration Director General Wu Ching-shan issued formal apologies in July 2014 for the confusion surrounding Taiwan’s new high school admissions system (Hung-kuo et al., 2014).
@@
In this sense, beyond their direct effect on choices and incentives, the TD mechanism’s deductions also act as a nudge, encouraging families to adopt more conservative school rankings.143143143While nudges are typically viewed as tools to guide individuals toward beneficial decisions—i.e., paternalistic nudges—in this case, they represent a form of nudging “for bad,” as discussed in Glaeser (2006); Thaler (2015); Schmidt and Engelen (2020).
@@ 13.4 Framing Institutions: How Descriptions Shape Their Adoption
Taiwan’s experience with its school choice mechanism illustrates how the framing of an institution can significantly influence both participant behavior and its acceptance by central planners, revealing the potential role of behavioral economics while exposing the limitations of traditional mechanism design in some practical contexts. In this abstract framework, a mechanism consists of a message space for each agent and an outcome function that maps message profiles to outcomes. It assumes that neither the process by which this mapping is derived nor how it is communicated affects participant behavior or the mechanism’s appeal to central planners. However, Taiwan’s case suggests that these assumptions may not always hold in real-world settings.144144144This observation aligns with the recent work of Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) on mechanism descriptions that make strategy-proofness easier to observe. Given a direct mechanism $\phi$, an individual $i$, and the reported types (preferences, in our context) of all other individuals, the menu for $i$ under $\phi$ is constructed by applying the mechanism across all possible types for $i$. Hammond (1979) shows that a direct mechanism is strategy-proof if and only if each agent always receives one of their best outcomes from their menu. Focusing on strategy-proof mechanisms, Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) suggest that the strategy-proofness of a mechanism can be made evident to participants through its menu descriptions, with experimental evidence indicating that truthful preference revelation increases under such descriptions. Their work highlights that, irrespective of the mechanism’s specific formulation, how it is communicated can significantly influence participant behavior.
@@
Consider the first formulation. As illustrated in Example 7, women are disadvantaged in the second batch of positions since their highest-merit-score members have already received assignments, while the highest-merit-score men remain. However, this disadvantage may not be obvious to the layperson.145145145Indeed, indicative of a conservative estimate of this oversight, an online experiment by Pathak et al. (2023) reveals that about 40% of subjects believe the order in which positions are allocated is irrelevant. In contrast, in the second formulation, it is clear that the benefits for women under this affirmative action policy are conditional, only activating if women do not receive 30 positions without affirmative action. Thus, the relative modesty of this policy is more apparent in the second formulation. Therefore, a decision maker seeking to adopt a more potent policy might be more inclined to choose the minimum guarantee choice rule when presented with its original formulation, perhaps overlooking its modesty, than if they were presented with the adjustment-based formulation.
As another example, consider the house allocation model with existing tenants, discussed in Section 5.6, and the equivalence presented in Theorem 2. Depending on which of the two formulations of the same mechanism is presented to decision makers, different reactions may be anticipated. One formulation clearly highlights the preferential treatment of newcomers over existing tenants, which may trigger moral concerns, while the other obscures this aspect through a core-based technocratic description, potentially avoiding such concerns.
Similarly, as we discussed in Section 6.10, when Boston Public Schools (BPS) was evaluating two strategy-proof mechanisms—DA and TTC—as potential replacements for the Boston mechanism, the Student Assignment Task Force formally recommended TTC. However, BPS leadership opted for DA due to various concerns related to the implicit “trading of priorities” carried out under TTC, a feature prominently highlighted in Abdulkadiroğlu and Sönmez (2003). Some of the concerns were that the trading shifted focus to priorities rather than the reasons they are awarded, created the possible impression that the mechanism could be manipulated (even though it could not), and led to a lack of transparency due to “behind-the-scenes mechanized trading.”

### v4

@@ 12.7 Malatya’s Journey to Global Leadership: Implementing Best Practices
The other two crucial factors are the capacity to conduct larger exchanges and, more importantly, the inclusion of compatible pairs in the system (see Figures 33 and 35). These are the areas in which the Malatya team truly excelled.
When our collaboration began in 2019, the Malatya team had already demonstrated the ability to perform five simultaneous liver transplants. Although the Liver Transplantation Institute at İnönü University was equipped to handle this capacity, they had the personnel to support up to seven. Following the world’s first 4-way exchange in July 2022, the team added two operating rooms to the existing ten at the Institute—all located on a single floor—enabling 6-way exchanges within its premises. As of September 2025, they had utilized their expanded capacity for 7-way exchanges twice, using additional operating rooms within the larger hospital complex.
However, while highly significant, this impressive logistical capacity was not the single most important driver of their extraordinary performance after the program’s pilot phase. Rather, it was the integration of “biological” incentive mechanisms—encouraging the inclusion of compatible pairs—that ultimately played the leading role in their success.
Following the highly successful media event at the end of July 2023, Yılmaz inquired about additional policies that could enhance the efficacy of the BBS–LPE System. Reminding him that our 4-way exchange in July 2022 had been made possible by a blood-type O donor paired with a blood-type A patient, we asked what other strategies might encourage such pairs to participate in liver exchange. He explained that there are several reasons why blood-type compatible pairs may benefit from participating in liver exchange.
@@
These scenarios have been utilized by the BBS–LPE System since its deployment. However, three additional factors could further enhance its efficacy.
Firstly, liver transplants using blood-type identical grafts are generally better for patients than those using blood-type compatible but non-identical grafts (Gordon et al., 1986; Koukoutsis et al., 2007). Although the patient may not have antibodies against the transplanted liver, the liver itself can temporarily carry antibodies against the patient’s body. This phenomenon, known as passenger lymphocyte syndrome, often leads to hemolytic anemia in recipients of blood-type non-identical grafts. While this condition typically resolves on its own, it can occasionally result in severe complications, including death. In fact, in medical practice, transplanting a blood-type compatible but non-identical graft is referred to as minor ABO incompatibility (Bruijn et al., 2017). This creates an additional “biological” incentive to enhance the efficacy of the BBS–LPE System.
@@
In addition to integrating these “biological” incentive mechanisms into the BBS–LPE System, the Malatya team has also allowed for “altruistic” participation from compatible pairs, even when there is no clear benefit for the pair, provided that the patient receives an equally suitable graft without delay.
The adoption of these policies since August 2023 has been the primary driver behind the exceptional outcomes of the BBS–LPE program following its pilot phase. Our hope is that these practices will evolve into established standards in liver exchange, enhancing its global efficacy by orders of magnitude.
@@ 13 Contributions, Insights and Directions on Behavioral Market Design
While I have primarily relied on theory as the formal apparatus for my pursuits in minimalist market design, many applications have revealed intriguing insights, questions, and puzzles in behavioral economics. In this section, I explore some of the most significant ones.
@@ 13.1 Leveling the Playing Field in School Choice
To build a compelling case against this mechanism in favor of the DA or SC–TTC mechanisms, Yan Chen and I conducted a laboratory experiment involving all three (Chen and Sönmez, 2006).154154154Building on Chen and Sönmez (2006), a substantial body of literature has employed lab experiments to study preference manipulation in school choice mechanisms. For a comprehensive review of this literature and other behavioral studies in education market design, see Rees-Jones and Shorrer (2023). For a broader review of behavioral studies in market design, see Chen et al. (2021). In our experimental setting, both DA and SC–TTC outperformed the Boston mechanism.
@@
In our experiment, we deliberately withheld this key information, not disclosing the strategy-proofness of DA and SC–TTC to participants. Our goal was to explore whether these mechanisms could still outperform Boston without any external guidance. Both mechanisms did, with DA performing best. However, a substantial number of participants still submitted non-truthful preferences under both DA and SC–TTC—albeit less frequently than under the Boston mechanism.
Encouraged by these experimental results, we grew confident that replacing the Boston mechanism with either DA or SC–TTC could yield even greater benefits in real-world settings, especially when families are properly informed and guided.
@@
Fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. The set of Nash equilibrium outcomes of this game is equal to the set of outcomes that satisfy individual rationality, non-wastefulness, and no justified envy.155155155Equivalently, invoking Proposition 4, the set of Nash equilibrium outcomes of this game is equal to the set of stable outcomes of the induced many-to-one matching market.
@@
Fix a school choice problem. The outcome of the DA mechanism under truthful dominant strategies either Pareto dominates or is equal to any given Nash equilibrium outcome of the preference revelation game induced by the Boston mechanism.156156156Ergin and Sönmez (2006) show that the complete information assumption is key for both Theorem 16 and Corollary 1. The characterization result in Theorem 16 not only fails to extend to an incomplete information environment, but in such an environment a student may also prefer a Bayesian Nash equilibrium outcome of the Boston mechanism to the outcome of the DA mechanism under truthful dominant strategies.
During our first meeting with Boston Public Schools (BPS) officials in October 2003, the experimental evidence from Chen and Sönmez (2006) and the theoretical results from Ergin and Sönmez (2006) were instrumental in helping to persuade BPS officials that it might be in the city’s best interest to consider replacing the Boston mechanism with either the DA or SC–TTC mechanism (see Section 6.10).
Even though the reform process took close to two years for due diligence, it was clear from the beginning that both BPS leadership and the members of the Task Force formed to evaluate the city’s assignment process were supportive of a reform that would increase the transparency of the student assignment mechanism. While the decision between DA and SC–TTC remained uncertain for some time, it was clear that there was little interest in retaining the Boston mechanism. Everything was progressing just as I had hoped.
@@
Subsequent analysis of school choice data from BPS under the Boston mechanism confirmed the widespread presence of both highly strategic and relatively naive preference submissions (Abdulkadiroğlu et al., 2006). For example, among the latter group, 19% of families listed two consistently oversubscribed (i.e., popular) schools as their top two choices, leading to 27% of these children being unassigned.
@@
Partitioning the students into strategic and sincere groups, fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. There exists a Nash equilibrium outcome in this game that is weakly preferred by each strategic student to the outcome of the DA mechanism in truthful dominant strategies.157157157With the BPS data used in Pathak and Sönmez (2008), the Nash equilibrium outcome of the preference revelation game induced by the Boston mechanism is virtually unique under all partitions of students into strategic and sincere.
@@
We now turn to a sharply contrasting case study from Seattle, where the school assignment process remained opaque for nearly a decade, possibly to serve objectives of decision-makers that, by my definition, do not qualify as “legitimate.”
@@ 13.2 Opting for Obscurity: The Curious Case of Seattle School Choice
Up until then, all of us in the market design community believed that no jurisdiction worldwide had used the DA mechanism for student assignment prior to the publication of Abdulkadiroğlu and Sönmez (2003). Thanks largely to MacGregor’s efforts, we discovered that Seattle had transitioned from the Boston mechanism to the DA mechanism in 1999, through what was known as the Barnhart–Waldman (BW) amendment. This amendment—named after two school board members familiar with the U.S. entry-level medical job market’s use of DA—led to the change. However, the switch was poorly communicated, if at all, resulting in widespread confusion among both the families participating in the system and the officials responsible for explaining it.
In Pathak and Sönmez (2013), we identified several indicators that the BW amendment was not well understood, even though a strategy-proof mechanism should have enabled the school district to advise families that submitting truthful preferences was a safe strategy. Confusion about the BW amendment even surfaced during a 2001 deposition of the school board president in a court challenge to Seattle’s school choice plan.158158158Page 58 of the U.S. Court of Appeals for the Ninth Circuit, No. 01–35450, Parents Involved in Community Schools vs. Seattle School District No.1, 2001.
@@
This also raises another issue. In Section 13.1, we have seen that even with full transparency, some participants in Boston utilized the available information better than others. In Seattle, judging from the inconsistent advice provided by officials, such as school principals, it is likely that while some families understood that truthful preference revelation was in their best interest, others did not. These families surely gained an even greater advantage than members of the WZPG in Boston compared to less informed families. In general, inconsistent communication of various aspects of an institution can even be exploited by decision-makers to benefit some groups at the expense of others.
@@
In June 2009, the Seattle School Board proposed a new student assignment plan to transition from the DA mechanism back to the Boston mechanism. The district publicly justified this reform as part of its “New Student Assignment Plan,” explicitly stating that “one of the goals of the new assignment plan is to reduce transportation costs” (Seattle Public Schools, 2009).
This proposal nevertheless faced significant opposition from families. As Walkup (2009) documents, many of the concerns echoed those raised in Abdulkadiroğlu and Sönmez (2003) regarding the limitations of the Boston mechanism.
@@
Opposition was not limited to technical critiques. Contemporaneous public commentary captured the depth of frustration with the reform. For example, a widely circulated blog post titled “Ten+ Reasons Why the Seattle Public Schools Superintendent Should Be Fired With Cause” listed the new assignment plan itself as a reason for dismissal, citing its effects on overcrowding, under-enrollment, and re-segregation (Seattle Education, 2011).159159159Indeed, the superintendent was dismissed shortly thereafter. In March 2011, the school board voted to remove her following revelations of financial improprieties, with board members noting that key audit findings had not been disclosed to them (Samuels, 2011).
@@
Perplexed by this reversal—the only instance we were aware of at the time in which a jurisdiction had replaced a school choice mechanism with a more manipulable one—we corresponded with several committee members involved in the decision. In Pathak and Sönmez (2013), we reported the only plausible explanation we encountered: that the BW amendment (i.e., the DA mechanism) would increase transportation costs by encouraging student mobility, as parents could freely express their true preferences without risking a seat at their neighborhood schools. This interpretation is fully consistent with the district’s own stated rationale in 2009, which explicitly identified reducing transportation costs as a goal of the new assignment plan (Seattle Public Schools, 2009). If this was indeed the case, it likely became a concern only after the strategy-proofness of the city’s DA mechanism was brought to public attention, thanks to MacGregor’s efforts.
If reducing transportation costs was indeed a factor in Seattle’s reversal to the Boston mechanism, it would exemplify an objective that does not fully meet my definition of “legitimate.” Rather than directly addressing school choice, decision-makers may have sought instead to limit the mobility it introduced indirectly by reverting to the Boston mechanism. This rationale would also be consistent with the decision to keep the strategy-proofness of the school choice mechanism opaque for nearly a decade—from 1999 to 2008.
@@
Beyond Seattle’s experience, could the Boston mechanism be a preferred mechanism for decision-makers who favor neighborhood-based assignment policies but avoid directly advocating for them due to political costs?
@@ 13.3 Nudging School Choices: Taiwan’s Deduction Mechanism
Despite these relatively modest deductions, it is evident that—partly due to the mechanism’s explicit penalties for lower-ranked choices—TD mechanisms other than DA are all susceptible to preference manipulation. Thus, of all TD mechanisms, only DA satisfies the axiom of strategy-proofness. Similarly, none of the TD mechanisms, except for DA, satisfy the axiom of no justified envy.160160160It’s important to note that the school priority rankings induced by the effective scores under the TD mechanism are artifacts of the mechanism, not inherent to the problem. Therefore, no justified envy is defined based on the original student scores from the CAE.
@@
In response, Education Minister Chiang Wei-ling and K–12 Education Administration Director General Wu Ching-shan issued formal apologies in July 2014 for the confusion surrounding Taiwan’s new high school admissions system (Hung-kuo et al., 2014).
@@
In this sense, beyond their direct effect on choices and incentives, the TD mechanism’s deductions also act as a nudge, encouraging families to adopt more conservative school rankings.161161161While nudges are typically viewed as tools to guide individuals toward beneficial decisions—i.e., paternalistic nudges—in this case, they represent a form of nudging “for bad,” as discussed in Glaeser (2006); Thaler (2015); Schmidt and Engelen (2020).
@@ 13.4 Framing Institutions: How Descriptions Shape Their Adoption
Taiwan’s experience with its school choice mechanism illustrates how the framing of an institution can significantly influence both participant behavior and its acceptance by central planners, revealing the potential role of behavioral economics while exposing the limitations of traditional mechanism design in some practical contexts. In this abstract framework, a mechanism consists of a message space for each agent and an outcome function that maps message profiles to outcomes. It assumes that neither the process by which this mapping is derived nor how it is communicated affects participant behavior or the mechanism’s appeal to central planners. However, Taiwan’s case suggests that these assumptions may not always hold in real-world settings.162162162This observation aligns with the recent work of Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) on mechanism descriptions that make strategy-proofness easier to observe. Given a direct mechanism $\phi$, an individual $i$, and the reported types (preferences, in our context) of all other individuals, the menu for $i$ under $\phi$ is constructed by applying the mechanism across all possible types for $i$. Hammond (1979) shows that a direct mechanism is strategy-proof if and only if each agent always receives one of their best outcomes from their menu. Focusing on strategy-proof mechanisms, Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) suggest that the strategy-proofness of a mechanism can be made evident to participants through its menu descriptions, with experimental evidence indicating that truthful preference revelation increases under such descriptions. Their work highlights that, irrespective of the mechanism’s specific formulation, how it is communicated can significantly influence participant behavior.
@@
Consider the first formulation. As illustrated in Example 11 (and depicted in Figure 25), women are disadvantaged in the second batch of positions because their highest-merit-score members have already received assignments, while the highest-merit-score men remain. However, this disadvantage may not be obvious to the layperson.163163163Indeed, indicative of a conservative estimate of this oversight, an online experiment by Pathak et al. (2023) reveals that about 40% of subjects believe the order in which positions are allocated is irrelevant. In contrast, in the second formulation, it is clear that the benefits for women under this affirmative action policy are conditional, only activating if women do not receive 30 positions without affirmative action. Thus, the relative modesty of this policy is more apparent in the second formulation. Therefore, a decision-maker seeking to adopt a more potent policy might be more inclined to choose the minimum guarantee choice rule when presented with its original formulation, perhaps overlooking its modesty, than if they were presented with the adjustment-based formulation.
As another example, consider the house allocation model with existing tenants, discussed in Section 5.6, and the equivalence presented in Theorem 2. Depending on which of the two formulations of the same mechanism is presented to decision-makers, different reactions may be anticipated. One formulation clearly highlights the preferential treatment of newcomers over existing tenants, which may trigger moral concerns, while the other obscures this aspect through a core-based technocratic description, potentially avoiding such concerns.
Similarly, as we discussed in Section 6.10, when Boston Public Schools (BPS) was evaluating two strategy-proof mechanisms—DA and SC–TTC—as potential replacements for the Boston mechanism, the Student Assignment Task Force formally recommended SC–TTC. However, BPS leadership opted for DA due to various concerns related to the implicit “trading of priorities” carried out under SC–TTC, a feature prominently highlighted in Abdulkadiroğlu and Sönmez (2003). Some of the concerns were that the trading shifted focus to priorities rather than the reasons they are awarded, created the possible impression that the mechanism could be manipulated (even though it could not), and led to a lack of transparency due to “behind-the-scenes mechanized trading.”